# Price formation without fuel costs: the interaction of elastic demand with storage bidding

### v1

@@ Appendix C Additional figures
In this section we show in a simplified case for a single year how prices duration curves are built up (Fig. S1), a sensitivity analysis for how price duration curves change if some dispatchable conventional is retained in the market (Fig. S2) or when some reserve storage discharging capacity is procured out-of-market (Fig. S3). We also show the heatmaps of average monthly baseload prices for different demand modeling scenarios with perfect and myopic foresight on 35 unseen test weather years (Fig. S4). Load duration curves corresponding to the scenarios shown in Figs. 3 and 6 respectively are shown in Figs. S5 and S6. In Fig. S7, we show the revenues of different system components by price band, similar to Mallapragada et al. [3]. In Fig. S8, we illustrate the operation of the electricity system as balance time series for different demand modelling and operational foresight cases. Finally, Fig. S9 shows the relation between hourly marginal storage values and the average annual capacity factor anomaly of wind and solar.
@@
Figure S1: Price duration and marginal storage value duration curves for a single year (2020) with inelastic demand up to a VOLL of 2000 €/MWh. Case a)/d) only has VRE (wind and solar) and unconstrained hydrogen storage without energy capacity cost. Since the storage itself does not need to recover any investment costs through price arbitrage, we see a uniform hydrogen price. This translates to four price levels in the electricity price duration curve; for VRE, VOLL, storage charging and discharging. Case b)/e) includes energy capacity costs for the hydrogen storage, creating two price levels for hydrogen such that the storage energy capacity can earn a revenue to recover its costs. This carries over to the electricity price duration curve, where there are now six price levels; two levels each for storage charging and discharging. Case c)/f) additional includes battery storage, which leads to a more complex price duration curve with the previous six major price levels and small steps at their boundary, resulting from the battery’s arbitrage operations.
@@
Figure S3: Sensitivity of price duration curves to reserve capacity in LT and ST model with 20 weather years from 2001-2020. Reserve capacity is forced into the system by setting a minimum capacity for the hydrogen reconversion to electricity. VOLL is 2000 €/MWh. Linear elastic demand run with parameters $a=2000$ and $b=10$. Overall, our findings remain unchanged with reserve capacity procured outside the market. For the elastic case, we have price identity between LT and ST model. For the inelastic case, the prices rise drastically in the ST model compared to the LT model as the reserve capacity is increased. The forced additional dispatch capacity adds scarcity prices to the storage medium, which is then reflected in the electricity prices. As this setup is outside the model equilibrium, total welfare is lower. These price increases are not seen in the elastic case, where the prices remain stable because the demand willingness to pay sets the price rather than the storage bid.
Figure S4: Heatmap of average monthly baseload prices over 35 unseen weather years for different demand modeling scenarios with perfect and myopic foresight with heuristic storage bidding. Years between 1951 and 2020 not indexed were used for the corresponding LT model.
@@
Figure S7: Cost recovery of each component by price band. In the inelastic case, most of the revenue is made in the price band of 1000-2000 €/MWh, while in the elastic case the revenue is more evenly distributed across different price bands. In both cases, storage buys electricity in low-priced hours and sells it in high-priced hours.
@@
Appendix D Main figures for different countries
In this section, we replicate the paper’s main figures for different countries, namely Spain and the United Kingdom. Figs. S10 and S11 correspond to the price duration curves shown for Germany in Fig. 3. Figs. S12 and S13 replicate the distribution of average annual and monthly baseload prices shown for Germany in Figs. 4 and 5. The comparison of price duration curves and cost recovery factors in different short-term models shown in Figs. S14 and S15 for Spain and the United Kingdom correspond to the comparisons shown for Germany in Figs. 6 and 7.
@@ Appendix D Main figures for different countries
Appendix E Additional tables
The tables shown in this section collect a range of metrics for the different cases examined. These metrics include, among other aspects, statistics on prices, load served, system cost, energy mix and capacities built. The tables cover the full 70-year optimizations for Germany (Table S2), Spain (Table S4), and the United Kingdom (Table S6), as well as the model runs where the long-term model was run on 35 different weather years than the 35 weather years the subsequent short-term model was run on for Germany (Table S3), Spain (Table S5), and the United Kingdom (Table S7). Additionally, Table S8 provides statistics for the cases where reserve capacity was forced into the system in Fig. S3
@@ Appendix E Additional tables
Table S2: Metrics for 70-year optimizations for Germany. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a-35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a-35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S3: Metrics for 35-year LT plus 35-year ST optimizations for Germany. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S4: Metrics for 70-year optimizations for Spain. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a-35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a-35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S5: Metrics for 35-year LT plus 35-year ST optimizations for Spain. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S6: Metrics for 70-year optimizations for the United Kingdom. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a-35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a-35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S7: Metrics for 35-year LT plus 35-year ST optimizations for the United Kingdom. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S8: Metrics for 20-year LT optimizations (2001-2020) for Germany with forced reserve capacity (hydrogen fuel cell). VOLL with 2000 €/MWh. Linear-elastic demand with $a=2000$ and $b=10$.

### v2

@@ Appendix E Additional figures
In this section we show in a simplified case for a single year how prices duration curves are built up (Fig. S2), a sensitivity analysis for how price duration curves change if some dispatchable conventional is retained in the market (Fig. S3) or when some reserve storage discharging capacity is procured out-of-market (Fig. S4). We also show the heatmaps of average monthly baseload prices for different demand modelling scenarios with perfect and myopic foresight on 35 unseen test weather years (Fig. S5). Load duration curves corresponding to the scenarios shown in Figs. 3 and 6 respectively are shown in Figs. S6 and S7. In Fig. S8, we show in which time periods demand reductions of varying magnitude are located. In Fig. S9, we show the revenues of different system components by price band, similar to Mallapragada et al. [4]. In Fig. S10, we show the distribution of annual revenues for different system components. In Fig. S11, we illustrate the operation of the electricity system as balance time series for different demand modelling and operational foresight cases. Finally, Fig. S12 shows the relation between hourly marginal storage values and the average annual capacity factor anomaly of wind and solar.
@@
Figure S2: Price duration and marginal storage value duration curves for a single year (2020) with perfectly inelastic demand up to a VOLL of 2000 €/MWh. Case a)/d) only has VRE (wind and solar) and unconstrained hydrogen storage without energy capacity cost. Since the storage itself does not need to recover any investment costs through price arbitrage, we see a uniform hydrogen price. This translates to four price levels in the electricity price duration curve; for VRE, VOLL, storage charging and discharging. Case b)/e) includes energy capacity costs for the hydrogen storage, creating two price levels for hydrogen such that the storage energy capacity can earn a revenue to recover its costs. This carries over to the electricity price duration curve, where there are now six price levels; two levels each for storage charging and discharging. Case c)/f) additional includes battery storage, which leads to a more complex price duration curve with the previous six major price levels and small steps at their boundary, resulting from the battery’s arbitrage operations.
@@
Figure S4: Sensitivity of price duration curves to reserve capacity in LT and ST model with 20 weather years from 2001-2020. Reserve hydrogen storage discharging capacity is forced into the system by setting a minimum capacity for the hydrogen turbine to electricity. VOLL is 2000 €/MWh. Linear elastic demand run with parameters $a=2000$ and $b=10$. Overall, our findings remain unchanged with reserve capacity procured outside the market. For the elastic case, we have price identity between LT and ST model. For the perfectly inelastic case, the prices rise drastically in the ST model compared to the LT model as the reserve capacity is increased. The forced additional dispatch capacity adds scarcity prices to the storage medium, which is then reflected in the electricity prices. As this setup is outside the model equilibrium, total welfare is lower. These price increases are not seen in the elastic case, where the prices remain stable because the demand willingness to pay sets the price rather than the storage bid.
Figure S5: Heatmap of average monthly baseload prices over 35 unseen weather years for different demand modelling scenarios with perfect and myopic foresight with heuristic storage bidding. Years between 1951 and 2020 not indexed were used for the corresponding LT model.
@@
Figure S8: Time periods with demand reductions of 90%, 80%, 70% and higher, shown for different demand modelling scenarios.
@@
Figure S9: Cost recovery of each component by price band. In the perfectly inelastic case, most of the revenue is made in the price band of 1000-2000 €/MWh, while in the elastic case the revenue is more evenly distributed across different price bands. In both cases, storage buys electricity in low-priced hours and sells it in high-priced hours.
Figure S10: Distribution of annual revenues per technology normalised to the long-term average revenue. Revenues shown are derived from the 70-year LT models with VOLL and PWL-elastic demand. Annual revenues show less year-to-year variation with demand elasticity, similar to average annual baseload prices in Fig. 4. Wind and solar revenues benefit from a natural volume-price hedge: higher availability lowers prices while lower availability raises them, resulting in even tighter revenue clustering.
@@
Appendix F Main figures for different countries
In this section, we replicate the paper’s main figures for different countries, namely Spain and the United Kingdom. Figs. S13 and S14 correspond to the price duration curves shown for Germany in Fig. 3. Figs. S15 and S16 replicate the distribution of average annual and monthly baseload prices shown for Germany in Figs. 4 and 5. The comparison of price duration curves and cost recovery factors in different short-term models shown in Figs. S17 and S18 for Spain and the United Kingdom correspond to the comparisons shown for Germany in Figs. 6 and 7. Figs. S19 and S20 show the price and load duration curves for scenarios with cross-elasticity terms. Fig. S21 shows the distribution of annual revenues per technology.
@@ Appendix F Main figures for different countries
Figure S19: Variant of Fig. S1 for Spain showing price and load duration curves for scenarios with cross-elasticity terms.
Figure S20: Variant of Fig. S1 for the United Kingdom showing price and load duration curves for scenarios with cross-elasticity terms.
a) Spain
b) United Kingdom
Figure S21: Variant of Fig. S9 showing the distribution of annual revenues per technology for Spain and the United Kingdom.
Appendix G Additional tables
The tables shown in this section collect a range of metrics for the different cases examined. These metrics include, among other aspects, statistics on prices, load served, system cost, energy mix and capacities built. The tables cover the full 70-year optimisations for Germany (Table S4), Spain (Table S6), and the United Kingdom (Table S8), as well as the model runs where the long-term model was run on 35 different weather years than the 35 weather years the subsequent short-term model was run on for Germany (Table S5), Spain (Table S7), and the United Kingdom (Table S9). Additionally, Table S10 provides statistics for the cases where reserve capacity was forced into the system in Fig. S4
@@ Appendix G Additional tables
Table S4: Metrics for 70-year optimisations for Germany. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S5: Metrics for 35-year LT plus 35-year ST optimisations for Germany. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S6: Metrics for 70-year optimisations for Spain. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S7: Metrics for 35-year LT plus 35-year ST optimisations for Spain. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S8: Metrics for 70-year optimisations for the United Kingdom. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
VOLL, LT, 35a VOLL, ST, 35a-35a, PF VOLL, ST, 35a-35a, myopic 48/24 VOLL, ST, 35a-35a, myopic 96/48 VOLL, ST, 35a-35a, myopic 96/48, C+5% VOLL, ST, 35a-35a, myopic 96/48, C-5% PWL-elastic default, LT, 35a PWL-elastic default, ST, 35a-35a, PF PWL-elastic default, ST, 35a-35a, myopic 48/24 PWL-elastic default, ST, 35a-35a, myopic 96/48 PWL-elastic default, ST, 35a-35a, myopic 96/48, C+5% PWL-elastic default, ST, 35a-35a, myopic 96/48, C-5%
@@
Table S9: Metrics for 35-year LT plus 35-year ST optimisations for the United Kingdom. (PF = perfect foresight; myopic X/Y = myopic foresight with X hours of foresight and Y hours of overlap; VOLL = perfectly inelastic demand up to value of lost load of 2000 €/MWh; PWL-elastic = Piecewise-linear elastic; C±$\pm$±X= capacity perturbation of $\pm$±X%; LT = long-term model; ST = short-term model; STD = Standard Deviation)
@@
Table S10: Metrics for 20-year LT optimisations (2001-2020) for Germany with forced reserve capacity (hydrogen turbine). VOLL with 2000 €/MWh. Linear-elastic demand with $a=2000$ and $b=10$.